OpenStudy (anonymous):

1. Kevin is a server at a restaurant. On average, he can serve 3 tables per hour. Kevin earns between $2.33 and $3.01 in gratuity per table. Which is a reasonable amount of gratuity that Kevin earned in a 27 hour work week?

OpenStudy (anonymous):

\[3\times 2.33\times 27\leq x\leq 3\times 3.01\times 27\]